Question 1
1.1 the following information is provided:
c= 0.8
t= 0.15
m= 0.2

Calculate the value of the multiplier based on the provided information for the following economies:

1.1.1 closed economy with no government
1.1.2 closed economy with government and tax
1.1 .3 open economy with a government, tax and induced imports.
1
Expert's answer
2020-04-22T11:48:55-0400

1.1.1

Y=C+I=0.8+0.2=1

"MPC=\frac{C}{Y}=\frac{0.8}{1}=0.8"

MPS=1-MPC=1-0.8=0.

"M=\frac{1}{1-MPC}=\frac{1}{0.2}=5"

1.1.2

let G=0.5

"M=\frac{1}{(1-MPC)(1-t)}\times G=\frac{1}{(1-0.8)(1-0.15)}\times0.5=2.94"

1.1.3

let induced imports=-0.2, f

"M=\frac{1}{(1-c)(1-t)+f}=\frac{1}{(1-0.8)(1-0.15)-0.2}=2.7"